# Transcendence and normality of complex numbers via Hurwitz continued fractions

### v2

@@ 5. First application: Borel hierarchy of the set of Hurwitz normal numbers
The objective of this section is to analyse the complexity of the set of Hurwitz normal numbers with respect to the Borel hierarchy employed in descriptive set theory. This hierarchy provides insight into the complexity for determining whether a complex number is a Hurwitz normal number. The main step of this proof is to obtain dynamical information of the subshift $\overline{\mathsf{R}}$ (Theorem 1.3). Subsequently, we leverage a result from Airey, Jackson, Kwietniak, and Mance to establish Theorem 5.1 below, which implies Theorem 1.2.
@@ 5.1. Hurwitz normal numbers
In fact, for almost every $z\in\mathfrak{F}$ the equality (12) holds for every finite regular word $\mathbf{b}$. We refer to such numbers $z$ as Hurwitz normal numbers and we denote them by $\operatorname{Norm}(\mu_{\operatorname{H}})$.
@@ 5.3. Feeble specification
Let $\mathscr{A}$ be non-empty and at most countable set. We denote non-empty words $\mathbf{u}$ on $\mathscr{A}$ as $\mathbf{u}=(u_{1},u_{2},\ldots)$. In symbolic dynamics, it is customary to represent words without parentheses, i.e., $\mathbf{u}=u_{1}u_{2}\cdots$, however, we decided to keep them.
@@
    There exists some $b\in\mathbb{Z}[i]$ such that $\operatorname{Pm}(b)\geq 3$ and $\mathfrak{F}_{n+1}(\mathbf{a}b)=\mathfrak{F}$.
@@
By definition, $\overline{\mathsf{R}}$ is closed in $\mathscr{D}^{\mathbb{N}}$. Since factors (subwords) of regular words are also regular, $\sigma[\mathsf{R}]\subseteq\mathsf{R}$. If $\mathbf{a}\in\mathsf{R}$ and $b=3+3i$, then $b\mathbf{a}\in\mathsf{R}$ and $\sigma(b\mathbf{a})=\mathbf{a}$, so $\mathsf{R}\subseteq\sigma[\mathsf{R}]$. Define the set
@@
    If $\mathbf{u},\mathbf{v}\in\mathscr{G}$, then $\mathbf{u}\mathbf{v}\in\mathscr{G}$ by Lemma 5.3.i.
@@
    Let $\varepsilon>0$ and $N\in\mathbb{N}$ such that $1/N<\varepsilon$. Take $\mathbf{v}=(v_{1},\ldots,v_{m})\in\mathscr{L}(\overline{\mathsf{R}})$ with $m\geq N$. Then $\mathbf{v}\in\mathsf{R}(m)$, so $(v_{1},\ldots,v_{m-1})\in\mathsf{R}(m-1)$ and, by Lemma 5.3.ii, there exists $a\in\mathscr{D}$ satisfying
@@
It is clear that $\mathscr{I}\subset\mathbb{C}$ is a Borel subset. We will use the functions $\Lambda$ and $\overline{\Lambda}\colon\overline{\mathsf{R}}\to\mathbb{C}$ defined in Subsection 4.1. By definition $\overline{\Lambda}$ is continuous.
@@
$E_{n}(\mathbf{b},M,k)$ :=$:=\left\{z=[0;a_{1},a_{2},\ldots]:\left|\frac{1}{M}\sum_{j=0}^{M-1}\chi_{\mathcal{C}_{n}(\mathbf{b})}\left(T^{j}(z)\right)-\mu_{\operatorname{H}}(\mathcal{C}_{n}(\mathbf{a}))\right|\leq\varepsilon_{k}\right\},$
@@
Since different regular cylinders of the same level have disjoint interior, from (13) we get
@@
Take any $n\in\mathbb{N}$ and $\mathbf{b}\in\mathsf{R}(n)$. Let $d\in\mathscr{D}$ be such $\operatorname{Pm}(d)\geq 3$ and $c\in\mathscr{D}$ such that
@@
(see Lemma 7.1 below). Since $d\mathbf{b}c\in\mathsf{R}(n+2)$ by Lemma 5.3. i, we may choose $k\in\mathbb{N}$ so large that
$\varepsilon_{k}<\frac{\mu_{\operatorname{H}}(\mathcal{C}_{n+1}(d\mathbf{b}c))}{4}$
@@
The choice of $k$ and the definition of $E_{n+1}(\mathbf{b}c,M,k)$ guaranty the existence of some $r\in\{n+1,\ldots,M+n\}$ satisfying
@@
Since $\operatorname{Pm}(d)\geq 3$, we may use Proposition 3.2 and induction on $|\mathbf{b}|$ that $\mathfrak{F}_{r}(\mathbf{d}^{\prime}d\mathbf{b}c)=\mathfrak{F}_{r}(\mathbf{b}c)$ and $\mathfrak{F}_{r-1}(\mathbf{d}^{\prime}d\mathbf{b})=\mathfrak{F}_{n}(\mathbf{b})$. Therefore, the second contention is (15) is equivalent to $\iota\tau_{c}[\overline{\mathfrak{F}}_{n+1}(\mathbf{b}c)]\subseteq\mathfrak{F}^{\circ}_{n}(\mathbf{b})$, which is equivalent to $\overline{\mathcal{C}}_{n+1}(\mathbf{b}c)\subseteq\mathcal{C}^{\circ}_{n}(\mathbf{b})$. We conclude $T^{r-n-1}(z)\in\mathcal{C}^{\circ}_{n}(\mathbf{b})$ and (14) follows. ∎
@@ 6. Second application: construction of transcendental numbers
We conjecture that the condition $|a_{n}|\geq\sqrt{8}$ for all $n\in\mathbb{N}$ can be removed in Theorem 6.2.2. The second application of our work, Theorem 1.4, is a construction of a family of transcendental numbers overlooked by Theorem 6.2. This provides evidence supporting our conjecture. A key ingredient of the proof is Schmidt’s extension of his celebrated Subspace Theorem to number fields, we only state it for $\mathbb{Q}(i)$. For $m\in\mathbb{N}$ and $\mathbf{z}=(z_{1},\ldots,z_{m})\in\mathbb{C}^{m}$, we write $\|\mathbf{z}\|=\max\{|z_{1}|,\ldots,|z_{m}|\}$ and $\overline{\mathbf{z}}=(\overline{z_{1}},\ldots,\overline{z_{m}})$.
@@
Let $\zeta,\mathbf{B}$ be as in Theorem 1.4 and let $(q_{n})_{n\geq 1}$ correspond to $\zeta$ as in Equation (5). Then, there exists an $\varepsilon>0$ such that for all $n\in\mathbb{N}$ we have $\psi^{2u_{n}}\geq|q_{2w_{n}}{q_{2w_{n}+2u_{n}+2v_{n}}}|^{\varepsilon}$.
@@
When $\liminf_{n\to\infty}w_{n}<\infty$, the sequence $(-2,1+iB_{1},-2,1+iB_{2},\ldots)$ also has finite repetition exponent and its corresponding sequence $(W_{n})_{n\geq 1}$ has a constant subsequence. The transcendence of $\zeta$ then follows from Theorem 6.2. Suppose that $w_{n}\to\infty$ when $n\to\infty$. As noted in [8, p. 1013], we do not lose any generality by assuming $B_{w_{n}}\neq B_{t_{n}}$ for all $n\in\mathbb{N}$. By Example 1, we have $\mathbf{a}\in\mathsf{Ir}$ and the output $\mathbf{b}$ of Algorithm 4.3 is
@@
Therefore, since for all $n\in\mathbb{N}$ we have $\mathfrak{F}_{2n}(\operatorname{pref}(\mathbf{b},2n))=\mathfrak{F}$ by Lemma 7.1 below and $u_{n}$ is even, the periodic sequence $Z_{n}:=W_{n}^{\prime}\,U_{n}^{\prime}\,V_{n}^{\prime}\,U_{n}^{\prime}\,V_{n}^{\prime}\,U_{n}^{\prime}\cdots$ belongs to $\overline{\mathsf{R}}$ and $\zeta_{n}:=\overline{\Lambda}(Z_{n})$ is a quadratic irrational over $\mathbb{Q}(i)$. Let $(p_{n})_{n\geq 0}$, $(q_{n})_{n\geq 1}$ correspond to $\mathbf{b}$ as in (5). Given the polynomial
$P_{n}(X):=\begin{vmatrix}q_{2w_{n}-1}&q_{2t_{n}-1}\\ q_{2w_{n}}&q_{2t_{n}}\end{vmatrix}X^{2}-\left(\begin{vmatrix}q_{2w_{n}-1}&p_{2t_{n}-1}\\ q_{2w_{n}}&p_{2t_{n}}\end{vmatrix}+\begin{vmatrix}p_{2w_{n}-1}&q_{2t_{n}-1}\\ p_{2w_{n}}&q_{2t_{n}}\end{vmatrix}\right)X+\begin{vmatrix}p_{2w_{n}-1}&p_{2t_{n}-1}\\ p_{2w_{n}}&p_{2t_{n}}\end{vmatrix},$
one can use the standard proof of Lagrange’s theorem on periodic continued fractions, for example, [13, Theorem 4.2], to check that $\zeta_{n}$ is a solution of $P_{n}$. We are going to use $P_{n}$ to obtain three collections of linear forms and then we will apply Theorem 6.3 to get a contradiction. From $\zeta,\zeta_{n}\in\overline{\mathcal{C}}_{2tn+2u_{n}}(b_{1},\ldots,b_{2tn+2u_{n}})$ and Proposition 3.1.v, we obtain $|\zeta-\zeta_{n}|\leq 2|q_{2t_{n}+2u_{n}}|^{-2}$ (in this estimate we use the triangle inequality as in the proof of Equation (10)). If $\lambda$ is as in Proposition 3.1.vi, we can use parts ii and v of Proposition 3.1 to obtain
@@
Certainly, by Proposition 3.1. ii, v and Equation (5),
@@
In the last line, the first inequality uses Proposition 3.1.vii. The definition of $\mathscr{M}_{1,1}$, $\mathscr{M}_{1,2}$, $\mathscr{M}_{1,3}$, $\mathscr{M}_{1,4}$, tells us that for each $n\in\mathbb{N}$ we have
@@
By Equation (17), Proposition 3.1.ii, and the boundedness of $\mathbf{B}$, we conclude that
@@
As a consequence, by (20), for $n\in\mathcal{N}_{1}$ we have $x_{n,2}=x_{n,3}$, hence
@@
so $\|\mathbf{v}_{n}\|_{\infty}\leq 2|q_{2w_{n}}q_{2t_{n}}|$. Consider the linear forms $\mathscr{L}_{2,1}$, $\mathscr{L}_{2,2}$, $\mathscr{L}_{2,3}$ in the variable $\mathbf{X}=(X_{1},X_{2},X_{3})$ and the linear forms $\mathscr{M}_{2,1}$, $\mathscr{M}_{2,2}$, $\mathscr{M}_{2,3}$ in the variable $\mathbf{Y}=(Y_{1},Y_{2},Y_{3})$ given by
@@
By (19), (22), and parts ii and v of Proposition 3.1, there are some constants $\kappa_{5},\kappa_{6},\kappa_{7}>0$ depending on $\zeta$ such that for all $n\in\mathcal{N}_{1}$ we have
@@
$\leq\kappa_{5}\frac{|q_{2t_{n}}|}{|q_{2w_{n}}q^{2}_{2t_{n}+2u_{n}}|}\begin{vmatrix}q_{2w_{n}-1}&q_{2t_{n}-1}\zeta-p_{2t_{n}-1}\\ q_{2w_{n}}&q_{2t_{n}}\zeta-p_{2t_{n}}\end{vmatrix}|q_{2w_{n}}|$
$\leq\kappa_{5}\frac{|q_{2w_{n}}|}{|q_{2t_{n}+2u_{n}}|}\leq\kappa_{5}\frac{|q_{2t_{n}}|}{|q_{2t_{n}+2u_{n}}|}\leq\frac{\kappa_{6}}{|q_{2w_{n}}q_{2t_{n}}|^{\varepsilon}}\leq\frac{\kappa_{7}}{\|\mathbf{v}_{n}\|^{\varepsilon}}.$
The next to last inequality uses Lemma 6.4 and Proposition 3.1.vii. Hence, Theorem 6.3 implies the existence of some infinite set $\mathcal{N}_{2}\subseteq\mathcal{N}_{1}$ and some $\mathbf{r}=(r_{1},r_{2},r_{3})\in\mathbb{Z}[i]^{3}$, $\mathbf{r}\neq(0,0,0)$, such that
@@
7. Proof of the algorithm
In this section, we demonstrate the validity of Algorithm 4.3, thereby proving Lemma 4.1. We begin by presenting some useful observations of HCF and then we prove Algorithm 4.3.
7.1. Preliminary lemmas
In most of the proofs to come, we use implicitly the equations (2) and (3).

### v3

@@ 5. First application: Borel hierarchy of the set of Hurwitz normal numbers
This section aims to analyze the complexity of the set of Hurwitz normal numbers with respect to the Borel hierarchy employed in descriptive set theory. This hierarchy provides insight into the intricacy for determining whether a complex number is a Hurwitz normal number. The main step of this proof is to obtain dynamical information of the subshift $\overline{\mathsf{R}}$ (Theorem 1.3). Subsequently, we leverage a result from Airey, Jackson, Kwietniak, and Mance to establish Theorem 5.1 below, which implies Theorem 1.2.
@@ 5.1. Hurwitz normal numbers
In fact, for almost every $z\in\mathfrak{F}$ the equality (13) holds for every finite regular word $\mathbf{b}$. We refer to such numbers $z$ as Hurwitz normal numbers and denote them by $\operatorname{Norm}(\mu_{\operatorname{H}})$.
@@ 5.3. Feeble specification
Let $\mathscr{A}$ be non-empty and at most countable set. We denote non-empty words $\mathbf{u}$ on $\mathscr{A}$ as $\mathbf{u}=(u_{1},u_{2},\ldots)$. In symbolic dynamics, it is customary to represent words without parentheses, i.e., $\mathbf{u}=u_{1}u_{2}\cdots$, however, we keep them.
@@
    There exists some $b\in\mathscr{D}$ such that $\operatorname{Pm}(b)\geq 3$ and $\mathfrak{F}_{n+1}(\mathbf{a}b)=\mathfrak{F}$.
@@
By definition, $\overline{\mathsf{R}}$ is closed in $\mathscr{D}^{\mathbb{N}}$. Since factors (subwords) of regular words are also regular, $\sigma[\overline{\mathsf{R}}]\subseteq\overline{\mathsf{R}}$. If $\mathbf{a}\in\overline{\mathsf{R}}$ and $b=3+3i$, then $b\mathbf{a}\in\overline{\mathsf{R}}$ and $\sigma(b\mathbf{a})=\mathbf{a}$, so $\overline{\mathsf{R}}\subseteq\sigma[\overline{\mathsf{R}}]$. Define the set
@@
    If $\mathbf{u},\mathbf{v}\in\mathscr{G}$, then $\mathbf{u}\mathbf{v}\in\mathscr{G}$ by Lemma 5.3.i..
@@
    Let $\varepsilon>0$ and $N\in\mathbb{N}$ such that $1/N<\varepsilon$. Take $\mathbf{v}=(v_{1},\ldots,v_{m})\in\mathscr{L}(\overline{\mathsf{R}})$ with $m\geq N$. Then $\mathbf{v}\in\mathsf{R}(m)$, so $(v_{1},\ldots,v_{m-1})\in\mathsf{R}(m-1)$ and, by Lemma 5.3.ii., there exists $a\in\mathscr{D}$ satisfying
@@
It is clear that $\mathscr{I}\subset\mathbb{C}$ is a Borel subset. We will use the continuous functions $\Lambda$ and $\overline{\Lambda}\colon\overline{\mathsf{R}}\to\mathbb{C}$ defined in Subsection 4.1.
@@
$E_{n}(\mathbf{b},M,k)$ :=$:=\left\{z=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}:\left|\frac{1}{M}\sum_{j=0}^{M-1}\chi_{\mathcal{C}_{n}(\mathbf{b})}\left(T^{j}(z)\right)-\mu_{\operatorname{H}}(\mathcal{C}_{n}(\mathbf{a}))\right|\leq\varepsilon_{k}\right\},$
@@
Since different regular cylinders of the same level have disjoint interiors, from (14) we get
@@
Take any $n\in\mathbb{N}$ and $\mathbf{b}\in\mathsf{R}(n)$. Let $d\in\mathscr{D}$ be such that $\operatorname{Pm}(d)\geq 3$ and $c\in\mathscr{D}$ such that
@@
(see Lemma 7.1 below). Since $d\mathbf{b}c\in\mathsf{R}(n+2)$ by Lemma 5.3. i., we may choose $k\in\mathbb{N}$ so large that
$\varepsilon_{k}<\frac{\mu_{\operatorname{H}}(\mathcal{C}_{n+2}(d\mathbf{b}c))}{4}$
@@
The choice of $k$ and the definition of $E_{n+1}(\mathbf{b}c,M,k)$ guaranty the existence of some $r\in\{n+2,\ldots,M+n-1\}$ satisfying
@@
Since $\operatorname{Pm}(d)\geq 3$, we may use Proposition 3.2 to show that
$\mathfrak{F}_{r}(\mathbf{d}^{\prime}d\mathbf{b}c)=\mathfrak{F}_{r}(\mathbf{b}c)\;\text{ and }\;\mathfrak{F}_{r-1}(\mathbf{d}^{\prime}d\mathbf{b})=\mathfrak{F}_{n}(\mathbf{b}).$
Therefore, the second contention is (16) is equivalent to $\iota\tau_{c}[\overline{\mathfrak{F}}_{n+1}(\mathbf{b}c)]\subseteq\mathfrak{F}^{\circ}_{n}(\mathbf{b})$, which is equivalent to $\overline{\mathcal{C}}_{n+1}(\mathbf{b}c)\subseteq\mathcal{C}^{\circ}_{n}(\mathbf{b})$. We conclude $z\in\mathcal{C}^{\circ}_{n}(d_{1},\ldots,d_{r-1})$ so $T^{r-n-1}(z)\in\mathcal{C}^{\circ}_{n}(\mathbf{b})$ and (15) follows. ∎
@@ 6. Second application: construction of transcendental numbers
We conjecture that the condition $|a_{n}|\geq\sqrt{8}$ for all $n\in\mathbb{N}$ can be removed in Theorem 6.2.2.. The second application of our work, Theorem 1.4, is constructing a family of transcendental numbers overlooked by Theorem 6.2. This provides evidence supporting our conjecture. A key ingredient of the proof is Schmidt’s extension of his celebrated Subspace Theorem to number fields, we only state it for $\mathbb{Q}(i)$. For $m\in\mathbb{N}$ and $\mathbf{z}=(z_{1},\ldots,z_{m})\in\mathbb{C}^{m}$, we write $\|\mathbf{z}\|=\max\{|z_{1}|,\ldots,|z_{m}|\}$ and $\overline{\mathbf{z}}=(\overline{z_{1}},\ldots,\overline{z_{m}})$.
@@
Let $\zeta,\mathbf{B}$ be as in Theorem 1.4 and let $(q_{n})_{n\geq 1}$ correspond to $\zeta$ as in Equation (6). Then, there exists an $\varepsilon>0$ such that for all $n\in\mathbb{N}$ we have $\psi^{2u_{n}}\geq|q_{2w_{n}}{q_{2t_{n}}}|^{\varepsilon}$.
@@
When $\liminf_{n\to\infty}w_{n}<\infty$, the sequence $(-2,1+iB_{1},-2,1+iB_{2},\ldots)$ also has finite repetition exponent and its corresponding sequence $(W_{n})_{n\geq 1}$ has a constant subsequence. The transcendence of $\zeta$ then follows from Theorem 6.2. Suppose that $w_{n}\to\infty$ when $n\to\infty$. As noted in [8, p. 1013], we do not lose any generality by assuming $B_{w_{n}}\neq B_{t_{n}}$ for all $n\in\mathbb{N}$. By Example 4.1, we have $\mathbf{a}\in\mathsf{Ir}$ and the output $\mathbf{b}$ of Procedure 4.3 is
@@
By Lemma 7.1 below and $\min_{j}|B_{j}|\geq 3$, we have $\mathfrak{F}_{2k}(\operatorname{pref}(\mathbf{b},2k))=\mathfrak{F}$ for all $k\in\mathbb{N}$. Then, since $u_{n}$ is even, the periodic sequence $Z_{n}:=W_{n}^{\prime}\,U_{n}^{\prime}\,V_{n}^{\prime}\,U_{n}^{\prime}\,V_{n}^{\prime}\,U_{n}^{\prime}\cdots$ belongs to $\overline{\mathsf{R}}$ and $\zeta_{n}:=\overline{\Lambda}(Z_{n})$ is a quadratic irrational over $\mathbb{Q}(i)$. Let $(p_{k})_{k\geq 0}$, $(q_{k})_{k\geq 0}$ correspond to $\mathbf{b}$ as in (6). One can use the standard proof of Lagrange’s theorem on periodic continued fractions (see, for example, [13, Theorem 4.2]) to check that $\zeta_{n}$ is a solution of the polynomial
$P_{n}(X):=\begin{vmatrix}q_{2w_{n}-1}&q_{2t_{n}-1}\\ q_{2w_{n}}&q_{2t_{n}}\end{vmatrix}X^{2}-\left(\begin{vmatrix}q_{2w_{n}-1}&p_{2t_{n}-1}\\ q_{2w_{n}}&p_{2t_{n}}\end{vmatrix}+\begin{vmatrix}p_{2w_{n}-1}&q_{2t_{n}-1}\\ p_{2w_{n}}&q_{2t_{n}}\end{vmatrix}\right)X+\begin{vmatrix}p_{2w_{n}-1}&p_{2t_{n}-1}\\ p_{2w_{n}}&p_{2t_{n}}\end{vmatrix}.$
We will use $P_{n}$ to obtain three collections of linear forms and then we will apply Theorem 6.3 to get a contradiction. From $\zeta,\zeta_{n}\in\overline{\mathcal{C}}_{2tn+2u_{n}}(b_{1},\ldots,b_{2tn+2u_{n}})$ and Proposition 3.1.(iv), we obtain $|\zeta-\zeta_{n}|\leq 2|q_{2t_{n}+2u_{n}}|^{-2}$ (in this estimate we use the triangle inequality as in the proof of Equation (11)). If $\lambda$ is as in Proposition 3.1.(v), we can use parts (ii) and (iv) of Proposition 3.1 to obtain
@@
Certainly, by Proposition 3.1. (ii), (iv) and Equation (6),
@@
In the last line, the first inequality uses Proposition 3.1.(vi). The definition of $\mathscr{M}_{1,1}$, $\mathscr{M}_{1,2}$, $\mathscr{M}_{1,3}$, $\mathscr{M}_{1,4}$, tells us that for each $n\in\mathbb{N}$ we have
@@
By Equation (18), Proposition 3.1.(ii), and the boundedness of $\mathbf{B}$, we conclude that
@@
As a consequence, by (21), for $n\in\mathcal{N}_{1}$ we have $x_{n,2}=x_{n,3}$ and
@@
so $\|\mathbf{v}_{n}\|\leq 2|q_{2w_{n}}q_{2t_{n}}|$. Consider the linear forms $\mathscr{L}_{2,1}$, $\mathscr{L}_{2,2}$, $\mathscr{L}_{2,3}$ in the variable $\mathbf{X}=(X_{1},X_{2},X_{3})$ and the linear forms $\mathscr{M}_{2,1}$, $\mathscr{M}_{2,2}$, $\mathscr{M}_{2,3}$ in the variable $\mathbf{Y}=(Y_{1},Y_{2},Y_{3})$ given by
@@
By (20), (23), and parts (ii) and (iv) of Proposition 3.1, there are some constants $\kappa_{5},\kappa_{6},\kappa_{7},\kappa_{8}>0$ depending on $\zeta$ such that for all $n\in\mathcal{N}_{1}$ we have
@@
$\leq\kappa_{5}\frac{|q_{2t_{n}}|}{|q_{2w_{n}}q^{2}_{2t_{n}+2u_{n}}|}\begin{vmatrix}q_{2w_{n}-1}&q_{2t_{n}-1}\zeta-p_{2t_{n}-1}\\ q_{2w_{n}}&q_{2t_{n}}\zeta-p_{2t_{n}}\end{vmatrix}|q_{2w_{n}}q_{2t_{n}}|$
$\leq\kappa_{6}\frac{|q_{2w_{n}}|}{|q_{2t_{n}+2u_{n}}|}\frac{|q_{2t_{n}}|}{|q_{2t_{n}+2u_{n}}|}$
$\leq\kappa_{6}\frac{|q_{2t_{n}}|}{|q_{2t_{n}+2u_{n}}|}\leq\frac{\kappa_{7}}{|q_{2w_{n}}q_{2t_{n}}|^{\varepsilon}}\leq\frac{\kappa_{8}}{\|\mathbf{v}_{n}\|^{\varepsilon}}.$
The next to last inequality uses Lemma 6.4 and Proposition 3.1.(vi). Hence, Theorem 6.3 implies the existence of some infinite set $\mathcal{N}_{2}\subseteq\mathcal{N}_{1}$ and some $\mathbf{r}=(r_{1},r_{2},r_{3})\in\mathbb{Z}[i]^{3}$, $\mathbf{r}\neq(0,0,0)$, such that
@@
7. Proof of Proposition 4.1
Recall that Proposition 4.1 says that for each $\mathbf{a}\in\mathsf{Ir}$ there is some $\mathbf{b}\in\overline{\mathsf{R}}$ such that $\overline{\Lambda}(\mathbf{b})=\Lambda(\mathbf{a})$. To prove it, we first need some auxiliary results.
7.1. Auxiliary lemmas
First, we gather some technical observations.